Help me!

11 A man pulls 100-Norabe shown. 20
Answer #1

here,

theta = 30 degree

force , F = 100 N

x = 5 m

the work done by the man , W = F * d

W = 100 * 5 = 500 J

the work done by the man is c. 500 J